### v1

# ST-USleepNet: A Spatial-Temporal Coupling Prominence Network for Multi-Channel Sleep Staging

Abstract
Sleep staging is critical for assessing sleep quality and diagnosing disorders. Recent advancements in artificial intelligence have driven the development of automated sleep staging models, which still face two significant challenges. 1) Simultaneously extracting prominent temporal and spatial sleep features from multi-channel raw signals, including characteristic sleep waveforms and salient spatial brain networks. 2) Capturing the spatial-temporal coupling patterns essential for accurate sleep staging. To address these challenges, we propose a novel framework named ST-USleepNet, comprising a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts raw signals into a spatial-temporal graph to model spatial-temporal couplings. The USleepNet utilizes a U-shaped structure originally designed for image segmentation. Similar to how image segmentation isolates significant targets, when applied to both raw sleep signals and ST module-generated graph data, USleepNet segments these inputs to extract prominent temporal and spatial sleep features simultaneously. Testing on three datasets demonstrates that ST-USleepNet outperforms existing baselines, and model visualizations confirm its efficacy in extracting prominent sleep features and temporal-spatial coupling patterns across various sleep stages. The code is available at: https://github.com/Majy-Yuji/ST-USleepNet.git.
1Saw Swee Hock School of Public Health, National University of Singapore, Singapore
2Institute of Automation, Chinese Academy of Sciences, Beijing, China
3Institute of Data Science, National University of Singapore, Singapore
1 Introduction
Sleep staging is essential for assessing sleep quality and diagnosing sleep disorders. Traditionally, sleep staging relies on the analysis of polysomnography (PSG) data, which includes mainly electroencephalogram (EEG) and other physiological signals such as electrooculogram (EOG), electromyogram (EMG), and electrocardiogram (ECG). These signals are divided into 30-second epochs, which are manually classified by experts into five stages—Wake, N1, N2, N3, and REM—according to the American Academy of Sleep Medicine (AASM) standards [1]. However, manual sleep staging is a labor-intensive process, prone to variability and human error, and lacks scalability.
To enhance the efficiency and accuracy of sleep staging, researchers have turned to artificial intelligence to develop automated sleep staging models [2, 3, 4]. Despite considerable advancements, there are also two unresolved challenges in this field.
Figure 1: (a) and (b) display characteristic sleep waveforms from different stages: (a) features spindle waves from N2, and (b) delta waves from N3. (c) presents brain networks and salient subnetworks during N2 and N3, with salient subnetworks formed by retaining only the green nodes. (d) illustrates different types of spatial-temporal relationships. Taking the green node as an example, its temporal correlations at different time points are shown by red arrows, its spatial dependence with other nodes at the same time by black arrows, and its spatial-temporal coupling with other nodes at different times by blue arrows.
How to simultaneously capture the characteristic sleep waveforms and salient spatial brain networks from multi-channel raw signals? Different sleep stages exhibit prominent temporal and spatial features, mainly including the characteristic sleep waveforms and salient spatial brain networks as outlined by AASM sleep standards [1]. For instance, Figure 1 illustrates the prominent sleep features of the N2 and N3 sleep stages. Figures (a) and (b) display the characteristic sleep waveforms over time, with the N2 stage marked by spindles and the N3 stage by delta waves. Figure 1 (c) shows the brain networks during the N2 and N3 stages, highlighting the subnetworks—salient spatial brain networks extracted by retaining only some nodes (the green nodes in the figure). It is evident that the subnetworks during the N2 and N3 stages are significantly different; during the lighter sleep of N2, brain activity is more pronounced, resulting in denser and more extensive network connections [5]. Given the importance of extracting prominent sleep features, SalientSleepNet [6] effectively extracts salient waveforms from single-channel signals and has achieved notable results. However, the work on single-channel signals cannot be directly extended to multi-channel signals. That is because multi-channel signals not only exhibit characteristic sleep waveforms across each single channel, as shown in Figures 1 (a) and (b) with characteristic sleep waveforms simultaneously in two channels, but also, as the signals interact, they form different functional connections at various sleep stages [7], thereby creating complex spatial patterns. Consequently, simultaneously capturing both multi-channel characteristic sleep waveforms and salient spatial brain networks is crucial for sleep staging and presents a significant challenge.
How to capture the spatial-temporal coupling patterns in sleep staging? Multiple brain regions exhibit complex yet crucial spatial-temporal coupling patterns as they evolve over time [8]. Figure 1 (c) illustrates the spatial-temporal relationships within the brain network over time. In the figure, each circle represents a PSG data channel, corresponding to electrode placement on the brain; each arrow signifies a connection within the brain network over time, with different arrows indicating different types of relations. Specifically, red arrows indicate temporal correlations within the same channel at different time points, while black arrows represent spatial dependence at the same time point across different channels. Given the complexity of brain network connections, there is also a spatial-temporal coupling feature between different channels at different times [9], indicated by blue arrows in the figure. Most existing models can only capture temporal correlations or spatial dependences. For instance, some CNN/RNN-based models [10, 11, 12, 13] can capture temporal correlations, while some GNN-based models [14, 15] can capture spatial dependences. Therefore, capturing these spatial-temporal coupling features for sleep staging presents a challenge.
To address the above challenges, we propose ST-USleepNet, a raw-data-based novel spatial-temporal coupling prominence framework for sleep staging. It consists of two main components: a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts multi-channel raw signals into a spatial-temporal graph, allowing for the modeling and capturing of spatial-temporal couplings. On the other hand, the USleepNet, which leverages a U-shaped structure originally developed for image segmentation [16], segments the prominent temporal and spatial sleep features of signals data and the graph constructed by the ST like how image segmentation isolates the main target of an image. Therefore, the USleepNet has two streams: for the temporal stream, we utilize a fully convolutional U-Net to segment characteristic sleep waveforms; for the spatial stream, we employ a graph convolutional U-Net to isolate salient spatial brain networks. The key contributions of this work include the following:
  • We develop a USleepNet, comprising a temporal prominence network and a spatial prominence network to simultaneously extract multi-channel characteristic sleep waveforms and salient spatial brain networks within each sleep stage.
  • We design a spatial-temporal graph construction module to model the spatial-temporal coupling patterns within a sleep segment.
  • Experimental results indicate that our ST-USleepNet model reaches state-of-the-art performance across three sleep staging datasets. Model visualizations reveal that our method can effectively extract prominent sleep features within different sleep stages.
2 Related Work
Initially, traditional machine learning methods such as Random Forest (RF) [17] and Support Vector Machine (SVM) [18] [19] were employed for sleep staging. However, most of these models heavily rely on expert knowledge, requiring the extraction of hand-crafted features before undertaking the sleep staging task, making them more time-consuming and expensive. To overcome this, researchers have started leveraging deep learning to bypass the need for manual feature engineering, primarily convolutional neural networks (CNNs) [11] [20] [12] [13] and Recurrent Neural Networks (RNNs) [21], to automatically extract high-dimensional deep features. However, such methods cannot be trained end-to-end.
Therefore, recent efforts have increasingly focused on refining end-to-end architectures that can efficiently process temporal correlations in sleep data without the preliminary steps required. Researchers mainly use CNNs [3] [6], RNNs [10], or a combination of both [2] [4], or Transformer [22] to model the temporal information for sleep staging. Perslev et al. [3] introduced the U-time model, which uses a CNN-based Unet to model single-channel signals. Jia et al. [6] proposed SalientSleepNet, which uses a $U^{2}$-structured CNN Unet combined with an attention mechanism to identify typical waveforms in single sleep signals. Phan et al. [22] used a transformer model and utilized attention scores to visualize sleep-related temporal features and the influence of different epochs. However, these methods mainly focus on capturing abstract temporal features instead of the Characteristic sleep waveforms across multiple channels, leading to limited accuracy.
However, the performance of temporal models suffers when modeling multi-channel signals because they ignore the spatial relations among brain regions, which are critically important [23]. Therefore, researchers have used Graph Neural Networks (GNNs) and Transformer models in sleep staging to capture spatial features. Jia et al. proposed GraphSleepNet [14], the first to apply a spatial-temporal graph convolution network (ST-GCN) to the sleep staging task. Subsequently, researchers introduced more graph-based methods. For example, Jia et al. proposed MSTGCN [20] and Ji et al. proposed JK-STGCN [15]. Both methods successfully modeled the spatial characteristics of the brain. Liu et al. proposed the BSTT [24] model, which uses a spatial-temporal transformer to capture spatial brain features. However, these methods mainly focus on capturing global spatial features instead of the salient spatial brain networks and ignore the spatial-temporal coupling patterns.
To address the limitation in existing methods and simultaneously capture the characteristic sleep waveforms and salient spatial brain networks from multi-channel raw signals, and capture the spatial-temporal coupling patterns within different sleep stages, we introduce ST-USleepNet for sleep staging.
3 Preliminary
In this study, the input is a segment of multi-channel physiological signals $s^{C\times f_{s}T}$, where $C$ represents the number of channels, $f_{s}$ denotes the sampling frequency, and $T$ indicates the duration of each segment. In sleep staging, it is common to have $T=30\,\text{s}$. The sleep staging problem is defined as the task of training a neural network $F$, to map a segment of multi-channel physiological signal data $s^{C\times f_{s}T}$, to a specific sleep stage $\hat{y}$. According to the AASM standard [1], each $\hat{y}\in{0,1,2,3,4}$ is matched to five sleep stages Wake, N1, N2, N3, and REM.
4 Methodology
Figure 2: The figure illustrates the overall framework in the central gray box while surrounding elements detail network layers and provide a simple example. There are two channels of one segment physiological signal in the upper left, represented in green and blue, as an example input. If the segment is viewed as three time intervals, then the input signals will be cut into 6 slices. The spatial-temporal graph construction module then forms a six-node graph and inputs it into encoders, where temporal and graph encoders downsample the feature of nodes (the initial feature is the raw slice signal) and graph respectively, yielding a subgraph with essential features. Taking this subgraph as input, the decoder includes a spatial decoder that restores the graph to a supergraph, and a temporal encoder that restores the graph’s feature dimensions, ultimately producing a feature graph identical in dimension to the original spatial-temporal graph. By concatenating the feature matrix of this graph according to the time and channel information, point-wise and segment-wise labels are generated.
The framework of ST-USleepNet is illustrated in Figure 2 and it consists of two primary modules. Initially, a graph construction module is developed to capture the spatial-temporal coupling patterns, converting raw signals into a spatial-temporal graph. Subsequently, we introduce USleepNet, which is designed to extract characteristic sleep waveforms and identify spatial salient brain networks from these signals. USleepNet is composed of two intertwined streams: the temporal stream, also called the temporal prominence network, represented by yellow blocks in Figure 2, and the spatial stream, also called the spatial prominence network, depicted by orange blocks. Each stream comprises $l$ encoder and $l$ decoder blocks, with $l=2$ exemplified in the figure. Additionally, the temporal stream integrates a segment classifier. To effectively combine the two streams, a fusion strategy is implemented.
4.1 Spatial-Temporal Graph Construction
The spatial-temporal graph construction module aims to transform the raw signals into a spatial-temporal graph. Given a multi-channel physiological signal data $s^{C\times f_{s}T}$, to fully capture the spatial-temporal coupling patterns within the segments, we cut the data $s^{C\times f_{s}T}$ into $n$ slices, resulting in $X^{nC\times f_{s}t}$. We use the raw signals $X$ directly as the initial features of the graph, then we have $G_{s}=(V,A^{nC\times nC},X^{nC\times f_{s}t})$, where $V$ represents the set of nodes, $A$ is the adjacency matrix, and $X$ is the features of nodes. The nodes of the graph $G_{s}$ are constituted by various slice data: node $N_{c_{i}t_{j}}$ represents a channel $c_{i}$, where $i\in\{k\in\mathbb{Z}\mid 0\leq k<C\}$, at a certain time point $t_{j}$, where $j\in\{k\in\mathbb{Z}\mid 0\leq k<n\}$. The feature corresponding to node $V_{c_{i}t_{j}}$ in $G_{s}$ is the raw signal of the respective slice ${X_{ij}}^{1\times f_{s}t}$. Next, we will discuss the construction of the adjacency matrix $A^{nC\times nC}$.
The adjacency matrix $A$ is derived from the multiplication of a similarity matrix with several weighted masks. First, an adjacency matrix is established based on the similarity between different features. This is because highly similar signals often indicate that they share similar or identical functional states. Cosine similarity is commonly used to quantitatively measure the similarity between two physiological signals:
$sim\_cos(X_{i},X_{j})=\frac{X_{i}\cdot X_{j}}{\sqrt{\sum_{k=1}^{n}X_{ik}^{2}}\sqrt{\sum_{k=1}^{n}X_{jk}^{2}}}$ (1)
where $X_{i}$ and $X_{j}$ represent features of two different nodes, respectively. However, not all nodes need to be assigned a weight, and establishing weighted edges between features with minimal similarity could negatively impact model training. Thus, a threshold $\delta$ is set, where an edge is only established between two nodes if their feature similarity exceeds $\delta$. Consequently, the similarity matrix $S$ can be established as follows:
$s_{ij}=\begin{cases}sim\_cos(X_{i},X_{j})&\text{if }sim\_cos(X_{i},X_{j})>\delta\\ 0&\text{otherwise}\end{cases}$ (2)
Next, we add some prior knowledge. One is to add the temporal information: slices closer in time should have a higher similarity. Therefore, a temporal weighted mask $M_{t}$ based on the time interval between two slices is constructed. Given a hyperparameter $\delta_{t}\in(0,1)$, the temporal weight ${m_{t}}_{ij}$ between the feature $X_{c_{i}}{t_{i}}$ and $X_{c_{j}}{t_{j}}$ can be expressed as:
${m_{t}}_{ij}=\delta_{t}^{|t_{i}-t_{j}|}$ (3)
Similarly, it is also necessary to consider the positional features of channels between slices. The similarity between slice signals from different channels should be lower than that between slices from the same channel. Therefore, define hyperparameters $\delta_{p}\in(0,1)$ to penalize the edge weights of features from different channels. Thus, the positional weight ${m_{p}}_{ij}$ between the feature $X_{c_{i}}{t_{i}}$ and $X_{c_{j}}{t_{j}}$ can be expressed as:
${m_{p}}_{ij}=\begin{cases}\delta_{p}&\text{if }c_{i}\neq c_{j}\\ 1&\text{otherwise}\end{cases}$ (4)
Therefore, the adjacency matrix $A$ for $G_{s}$ can be represented as:
$A=S\cdot M_{t}\cdot M_{p}$ (5)
According to the above rules, a segment of the original multi-channel physiological signal data $s^{C\times f_{s}T}$ can be transformed into a spatial-temporal graph $G_{s}=(V,A^{nC\times nC},X^{nC\times f_{s}t})$.
4.2 U-shaped Sleep Network
The spatial-temporal graph $G_{s}$ constructed from the raw signals is then input into the U-shaped sleep network (USleepNet). There are two interwoven networks, the temporal prominence network for the temporal stream and the spatial prominence network for the spatial stream in USleepNet.
4.2.1 Temporal Prominence Network
We propose the temporal prominence network for the temporal stream. This network consists of three logical submodules: the temporal encoder block $T_{en}$, the temporal decoder block $T_{de}$, and a segment classifier $T_{c}$. $T_{en}$ takes the feature matrix $X$ of the graph $G_{s}$ as input, extracts temporal features at different scales, and outputs a compressed feature representation. Then, $T_{de}$ continues to learn from the compressed feature representation provided by $T_{en}$ to obtain a point-wise label for each point. Last, $T_{c}$ integrates the point-wise labels into a segment-wise label.
Temporal Encoder Block
The temporal encoder block $T_{en}$ consists of two 2D convolution layers. The number of channels in the second convolution layer is greater than that in the first layer since deeper networks tend to capture more complex features. Each convolution uses the same padding, followed by batch normalization and dropout. The kernel size remains consistent within the two convolution operations of each $T_{en}$, while it may vary across different ${T}{en}$. After these two convolution operations, max pooling is performed to reduce the feature dimensions.
Temporal Decoder Block
The temporal decoder block $T_{de}$ consists of two transposed-2D convolution blocks. Each $T_{de}$ corresponds one-to-one with a $T_{en}$, using inverse operations to gradually recover the features. First, the compressed feature representation is unpooled. Next, to retain more detail and merge features of different scales, $T_{de}$ concatenates the unpooled vector with the output of the corresponding $T_{en}$. Next, the concatenated vector undergoes two convolution operations sequentially. $T_{de}$ shares some similarities with $T_{en}$, such as the kernel size remaining consistent within the two convolution operations of each $T_{de}$ but may vary across different $T_{de}$; and each CNN layer is followed by batch normalization and dropout layers.
Segment Classifier
The segment classifier $T_{c}$ is a module to get the segment-wise label, primarily composed of a fully connected layer. The input to ${T}_{c}$ is an embedding $H^{ch\times nC\times f_{s}t}$, where $ch$ is the channel number. First, $H$ is reshaped into ${\hat{H}}^{ch\times C\times nf_{s}t}$. Next, a softmax function is used to obtain the classification for each time point of each physiological signal channel, considered as point-wise labels. Then, mean pooling is used to integrate the channel $h$ into one, and a fully connected layer is used to obtain the final label.
4.2.2 Spatial Prominence Network
For the spatial stream, we propose the spatial prominence network. This network follows an encoder-decoder structure and consists of two logical submodules, the spatial encoder block $S_{en}$ and the spatial decoder block $S_{de}$. $S_{en}$ takes the spatial-temporal graph $G_{s}$ as input and outputs a compressed feature subgraph, while ${S}{de}$ takes the subgraph and recovers a supergraph containing the learned spatial relationships.
Spatial Encoder Block $S_{en}$ is to extract spatial features between multiple channels and sample some important channels to form a new subgraph for high-level feature encoding. It consists of two consecutive GCN layers and a graph pooling layer.
The input to $S_{en}$ consists of the adjacency matrix $A^{v\times v}$ and the feature matrix $X^{v\times h}$ of the spatial-temporal graph $G_{s}$, where $v$ represents the number of nodes and $h$ represents the feature dimensions. The first graph convolution layer maps $X$ to a higher-dimensional space, and the second graph convolution layer maps $X$ back to its original dimensions. The two consecutive graph convolutions can be represented as follows:
$\mathbf{X}^{(2)}=Drop(\sigma(\mathbf{A}Drop(\sigma(\mathbf{A}\mathbf{X}\mathbf{W}_{1}^{h\times\hat{h}}))\mathbf{W}_{2}^{\hat{h}\times h})$ (6)
where $Drop$ means the dropout layder, and $\sigma$ is ELU, which is defined as:
$\sigma(x)=\text{ELU}(x)=\begin{cases}x&\text{if }x>0\\ \alpha(e^{x}-1)&\text{if }x\leq 0\end{cases}$ (7)
Then, a 2D graph pooling layer (2D gPool) is used to perform down-sampling on the graph data. This reduces the size of the graph while retaining information on important nodes. In 2D gPool, we rank the nodes based on importance and select k to generate a subgraph. The computation of gPool can be represented as follows: First, each node’s feature vector $X_{i}$ is projected using a trainable projection vector $P$ to calculate a scalar projection value as an importance score $ic$:
$ic^{v}=\frac{X^{v\times h}p^{h}}{|p^{h}|}$ (8)
where $v$ represents the number of nodes and $h$ represents the feature dimension. Then, sort the $ic$ and select k largest values:
$\text{idx}=\text{rank}(ic,k)$ (9)
$\tilde{ic}^{k}=sigmoid(ic(\text{idx})^{k})$ (10)
where $idx$ is the number of selected k nodes, $\tilde{ic}^{k}$ is the weight for the top k nodes. Select the feature vectors corresponding to the top k nodes and update the adjacency matrix $\tilde{A}$ and the feature matrix $\tilde{X}$:
$\tilde{X}^{v\times k}=X^{v\times h}(\text{idx},:)$ (11)
$\tilde{A}^{k\times k}=A^{v\times v}(\text{idx},\text{idx})A^{v\times v}(\text{idx},\text{idx})$ (12)
$\tilde{X}^{k\times h}=\tilde{X}^{v\times h}\odot(\tilde{ic}\mathbf{1}_{h}^{T})$ (13)
where $\odot$ denotes element-wise multiplication, and $\mathbf{1}_{h}$ represents an h-dimensional constant vector of 1 used to expand the dimensions of $\tilde{ic}$. In equation 12, the reason for multiplying the two extracted adjacency matrices is to avoid the nodes in the pooled graph becoming isolated as much as possible.
Spatial Decoder Block The Spatial Decoder Block $S_{de}$ is used to recover the compressed subgraph obtained by $S_{en}$ into a feature supergraph. It consists of four network layers: a convolution layer, a 2D graph unpooling layer (2D gUnpool), and two graph convolution layers.
$S_{de}$ takes the adjacency matrix $A$ and the feature matrix $X$ of a spatial-temporal graph $G_{s}$ as input. First, a CNN layer is used to extract local information from the feature matrix $X$ and reduce the number of channels in $X$. Next, the data is upsampled using a 2D gUnpool layer. The 2D gUnpool layer corresponds to a 2D gPool layer and uses the node selection information recorded during the downsampling process of the 2D gPool layer. It places the existing nodes and their feature matrices back into their original positions in the graph, while the features of other nodes are set to zero. To avoid significant information loss, a skip connection is used to add the adjacency matrix of the upsampled graph to the corresponding adjacency matrix before downsampling in $S_{de}$. Finally, similar to $S_{en}$, two consecutive graph convolution layers are used: the first maps the feature matrix to a hidden high-dimensional space, and the second restores it to its original dimensions. Through these four layers, the $S_{de}$ block ultimately recovers the compressed feature subgraph into a feature supergraph.
4.2.3 Spatial-Temporal Block Fusion
The USleepNet is composed of an alternating combination of the temporal prominence network and the spatial prominence network. USleepNet includes $l$ layers of spatial-temporal encoders ${ST}_{en}$ and $l$ layers of spatial-temporal decoders ${ST}_{de}$, with a segment classifier at the end. Specifically, ${ST}_{en}$ consists of a $T_{en}$ followed by an $S_{en}$, and ${ST}_{de}$ consists of an $S_{de}$ followed by a $T_{de}$. In the encoder phase, processing with $T_{en}$ first is essential because PSG data usually has high dimensions and contains a lot of noise. Convolution and pooling operations can reduce the dimensions while removing noise, and extracting more representative features. These lower-dimensional features can reduce the computational complexity of subsequent graph processing.
5 Experiments
5.1 Datasets
We evaluate the ST-USleepNet on three public datasets. In this part, We briefly describe each dataset and refer to the original papers for details.
ISRUC-S1 [25] contains PSG data from 100 patients (55 male and 45 female), with a total of 87,187 sleep segments, each lasting 30 seconds. The PSG data includes 10 channels of EEG, EOG, and EMG. ISRUC-S3 [25] contains PSG data from 10 patients (9 male and 1 female), with a total of 8,589 sleep segments. The PSG data includes 10 channels of EEG, EOG, and EMG. MASS-SS3 [26] contains PSG data from 62 healthy people (28 male and 34 female), with a total of 59,304 sleep segments. The PSG data includes 25 channels of EEG, EOG, EMG, and ECG.
5.2 Experiment Settings
To comprehensively evaluate the performance of ST-USleepNet, we select 9 models as baselines. These include classic works in sleep staging [27, 2, 4]; sleep networks based on U-shaped structures [3, 6]; spatial-temporal sleep networks [14, 24]; and the latest general multivariate time series model [28].
In the experiments, we divide the dataset into train, validation, and test sets, comprising 80%, 10%, and 10% of the total data, respectively. We select the best-performing model on the validation set and then test it on the test set. The evaluation metrics used are accuracy and F1 score.
All experiments are performed using NVIDIA A100 Tensor Core GPU. We configure the batch size to 64, utilize the Adam optimizer with a learning rate of 1e-3, and train the model for 80 epochs. Additional information is available in the appendix.
5.3 Experiment Results
We compare ST-USleepNet with baseline models on three public datasets. The results are shown in Figure 1, where the best results for each metric are highlighted in bold and the second best is underlined. Our model achieves state-of-the-art performance across all metrics and demonstrates significant improvements compared to existing methods. On the three datasets, ST-USleepNet shows improvements over existing methods in accuracy by 1.7%, 4.2%, and 0.9%, and in F1 score by 1.8%, 6%, and 3.3%, respectively. This underscores the importance of extracting prominent sleep features and capturing spatial-temporal coupling patterns for sleep staging.
Table 1: Performance comparison with baseline models.
Dataset ISRUC-S1 ISRUC-S3 MASS-SS3
Model Acc F1 Acc F1 Acc F1
RF 66.4 60.8 70.2 68.8 69.1 65.4
SVM 67.3 62.8 71.8 60.2 69.3 66.7
DeepSleepNet 73.0 69.1 75.7 73.2 84.8 79.8
TinySleepNet 75.3 73.9 76.1 74.5 82.4 77.8
U-time 71.8 67.1 75.5 71.9 85.0 79.3
GraphSleepNet 77.3 74.7 80.6 78.2 87.2 85.9
SalientSleepNet 80.6 80.0 80.3 78.4 84.9 82.8
BSTT 77.7 76.1 81.9 80.3 88.7 86.0
FC-STGNN 76.4 73.4 78.3 76.8 86.2 85.1
Ours 82.3 81.8 86.1 86.3 89.6 89.3
5.4 Ablation Study
To verify the effectiveness of the modules designed in ST-USleepNet, we design three variant models by removing the corresponding modules: 1) ST-USleepNet without temporal prominence network (-T). 2) ST-USleepNet without spatial prominence network (-S). 3) ST-USleepNet without spatial-temporal graph construction (-ST). In the ’-ST’ variant mode, we construct a spatial-only graph using the raw EEG signals in the whole segment based on the cosine similarity.
Figure 3: Results of ablation study on MASS-SS3 dataset.
Figure 3 shows the comparison results of the ablation experiments between the variant models and the original model on the MASS-SS3 dataset. It can be seen that removing either the temporal prominence network or the spatial prominence network adversely affects the model’s performance, and the impact is similar. This indicates that extracting characteristic sleep waveforms and salient spatial brain networks are both important. Additionally, removing the temporal-spatial graph construction module also affects model performance, suggesting that we should also focus on the spatial-temporal coupling patterns contained throughout the entire sleep signal, rather than just temporal or spatial features. Additionally, we explore the impact of important hyperparameters on the model’s performance. Detailed results are provided in the appendix.
5.5 Model Visualization
In the ST-USleepNet, each module can extract one kind of sleep feature: The temporal prominence Network can simultaneously capture characteristic sleep waveforms from multiple channels; The spatial prominence Network can capture the salient spatial Brain networks; The spatial-temporal graph construction module could capture the spatial-temporal coupling patterns. We show the visualizations for these three kinds of sleep features in this section.
Figure 4: Visualization results of extracted features.
5.5.1 Characteristic Sleep Waveforms
We visualize the model’s point-wise label output for each channel, demonstrating that the ST-USleepNet can simultaneously capture the characteristic sleep waveforms of various stages in multi-channel sleep data. As shown in Figure 4 (a), examples of characteristic sleep waveforms captured by the model for the five sleep stages are displayed. Each stage shows the results for two channels: for the W, N1, N2, and N3 periods, the F3 and C3 EEG channels are selected; for the REM period, the F3 EEG channel and the LOC EOG channel are chosen. It is evident that the model can simultaneously capture multi-channel wake period $\alpha$ waves; N1 period vertex waves, N2 period spindles, N3 period $\delta$ waves, and during the REM period, the EEG’s sawtooth waves and EOG’s rapid eye movement. In summary, our model effectively extracts characteristic sleep waveforms from multi-channel PSG data simultaneously.
5.5.2 Salient Spatial Brain Networks
We average the spatial-temporal graphs obtained from USleepNet into a single spatial graph and extract the salient brain networks according to the node of subgraphs output by the bottom block. The visualization results are in Figure 4 (b). In the figure, the left picture displays the complete brain network and several other nodes like EOG, while the right pictures show the extracted salient brain networks for each sleep stage. It is evident that the salient brain networks vary significantly across different sleep stages. More, specifically, during the wake stage, all brain regions are relatively active; in the N1 stage, as sleep onset begins, brain activity intensifies and eye movement reduces compared to the Wake stage; in the light sleep N2 stage, eye movement activity weakens further; in the N3 stage, deep sleep is entered, and brain network activity decreases; and during the REM stage, brain network activity is weaker, but eye movement is notably active. Thus, the ST-USleepNet can extract salient brain networks and possesses a certain degree of interpretability.
5.5.3 Spatial-Temporal Coupling Patterns
We visualize the spatial-temporal graph output by USleepNet as shown in Figure 4 (c). In the figure, to highlight the spatial-temporal coupling, the spatial layout at each time point is omitted in the drawing, and different physiological signal nodes are represented by distinct colors and arranged in a single column. It is evident that the ST-USleepNet effectively captures spatial-temporal coupling patterns, and the patterns reflected are consistent with the salient spatial brain network.
6 Conclusion
We propose ST-USleepNet, a novel spatial-temporal coupling prominence network for sleep staging. Our model can simultaneously capture characteristic sleep waveforms and salient spatial brain networks across multiple channels. Besides, by constructing the raw signals to graph, we capture the spatial-temporal coupling patterns for different stages. Experiments result shows that ST-USleepNet reaches the state-of-the-art performance. visualization results prove that our model has a certain degree of interpretability.
7 Acknowledgement
This research is supported by A*STAR, CISCO Systems (USA) Pte. Ltd, and National University of Singapore under its Cisco-NUS Accelerated Digital Economy Corporate Laboratory (Award I21001E0002) and the National University of Singapore President’s Graduate Fellowship.
Appendix
Figure S. 1: A 30-second segment of a typical polysomnography (PSG) recording, displaying multiple electrooculogram (EOG), electroencephalogram (EEG), electromyogram (EMG), and electrocardiogram (ECG) channels, sourced from the MASS-SS3 dataset. Human experts evaluate such segments and classify them into one of the sleep stages: Wake, Non-REM 1 (N1), Non-REM 2 (N2), Non-REM 3 (N3), or REM. ST-USleepNet considers all available channels for analysis.
Table S. 1: Characteristic sleep Waveforms in different sleep stages
Sleep Stage Label Description
Wake Wake Characterized by alpha waves (8-13 Hz) during relaxed wakefulness and beta waves (13-30 Hz) during active mental activity or alertness. Eye movements can include rapid movements or reading saccades, with occasional blinks.
Non-REM 1 N1 Dominated by theta waves (4-7 Hz). This light sleep stage often shows slow rolling eye movements as the transition from wakefulness to sleep begins. Vertex waves (sharp transients typically seen at the vertex of the scalp) are also characteristic. Some low-level muscle (EMG) activity persists.
Non-REM 2 N2 Marked by the presence of sleep spindles (short bursts of 13-16 Hz activity) and K-complexes (sharp negative deflections followed by a positive wave). Theta waves continue to dominate between these distinctive patterns.
Non-REM 3 N3 Characterized by high amplitude, slow delta waves (0.5-3 Hz). This stage, also known as slow-wave sleep (SWS), reflects deep sleep. Very low EMG activity is observed, and delta waves dominate the EEG.
REM REM Features low-amplitude, mixed-frequency EEG activity resembling wakefulness, with rapid eye movements (REMs). Theta waves and sawtooth waves (brief, rhythmic 2-6 Hz waves) are common. EMG activity is markedly reduced, often appearing as muscle atonia, while REMs are a defining characteristic.
Table S. 2: Sample number and sample propotion in each dataset.
Dataset Label Training Set Validation Set Test Set Total
Number(Proportion) Number(Proportion) Number(Proportion) Number(Proportion)
ISRUC-S1 Wake 16,658(23.9%) 2,010(23.1%) 2,132(24.5%) 20,800(23.9%)
N1 8,913(12.8%) 1,106(12.7%) 1,112(12.7%) 1674(12.8%)
N2 21,701(31.1%) 2,751(31.6%) 2,625(30.1%) 27,077(31.1%)
N3 13,704(19.7%) 1,725(19.8%) 1,703(19.5%) 17,168(19.7%)
REM 8,737(12.5%) 1,127(12.9%) 1,147(13.2%) 11,011(12.6%)
Total 69,749(100.00%) 8,719(100.00%) 8,719(100.00%) 87,187(100.00%)
ISRUC-S3 Wake 1,340(19.5%) 167(19.4%) 167(19.4%) 1,674(19.4%)
N1 973(14.2%) 122(14.2%) 122(14.2%) 1,674(14.2%)
N2 2,092(30.4%) 262(30.5%) 262(30.5%) 2,616(30.5%)
N3 1,614(23.4%) 201(23.4%) 201(23.4%) 2,016(23.4%)
REM 852(12.5%) 107(12.5%) 107(12.5%) 1,066(12.5%)
Total 6,871(100.00%) 859(100.00%) 859(100.00%) 8,589(100.00%)
MASS-SS3 Wake 5,030(10.6%) 644(10.8%) 667(11.3%) 6,341(10.7%)
N1 3,806(8.0%) 484(8.2%) 450(7.6%) 4,740(8.0%)
N2 23,738(50.0%) 2,979(50.0%) 2,970(50.0%) 29,687(50.0%)
N3 6,174(13.0%) 765(13.0%) 781(13.2%) 7,720(13.0%)
REM 8,696(18.4%) 1,058(17.8%) 1,062(17.9%) 10,816(18.3%)
Total 47,444(100%) 5,930(100%) 5,930(100%) 59,304(100%)
Table S. 3: Detailed information of channels used in each datset
  DataSet Signal type Label Description
  ISRUC-S1 EOG LOC-A2 Left eye movements.
ROC-A1 Right eye movements.
EEG F3-A2 Frontal, left hemisphere, referenced to the right earlobe.
C3-A2 Central, left hemisphere, referenced to the right earlobe.
O1-A2 Occipital, left hemisphere, referenced to the right earlobe.
F4-A1 Frontal, right hemisphere, referenced to the left earlobe.
C4-A1 Frontal, right hemisphere, referenced to the left earlobe.
O2-A1 Occipital, right hemisphere, referenced to the left earlobe.
EMG X1 Placed between the chin and the lower lip.
X2
  ISRUC-S3 EOG LOC-A2 Left eye movements.
ROC-A1 Right eye movements.
EEG F3-A2 Frontal, left hemisphere, referenced to the right earlobe.
C3-A2 Central, left hemisphere, referenced to the right earlobe.
O1-A2 Occipital, left hemisphere, referenced to the right earlobe.
F4-A1 Frontal, right hemisphere, referenced to the left earlobe.
C4-A1 Frontal, right hemisphere, referenced to the left earlobe.
O2-A1 Occipital, right hemisphere, referenced to the left earlobe.
EMG X1 Placed between the chin and the lower lip.
X2
  MASS-SS3 EEG Fp1 Frontal pole, left hemisphere.
Fp2 Frontal pole, right hemisphere.
Fz Frontal midline.
F3 Frontal, left hemisphere.
F4 Frontal, right hemisphere.
F7 Frontal, left hemisphere, closer to temporal.
F8 Frontal, right hemisphere, closer to temporal.
C3 Central, left hemisphere.
C4 Central, right hemisphere.
T3 Temporal, left hemisphere.
T4 Temporal, right hemisphere.
Pz Parietal midline.
P3 Parietal, left hemisphere.
P4 Parietal, right hemisphere.
T5 Temporal, left hemisphere, posterior.
T6 Temporal, right hemisphere, posterior.
Oz Occipital midline.
O1 Occipital, left hemisphere.
O2 Occipital, right hemisphere.
EOG EogL Left eye movements (Electrode near left eye).
EogR Right eye movements (Electrode near right eye).
EMG Emg1 Electromyography, placement varies.
Emg2 Electromyography, placement varies.
Emg3 Electromyography, placement varies.
ECG Ecg Electrocardiogram, measures heart activity.
Table S. 4: ST-USleepNet model topology. The layer dimensions listed below are based on the MASS-SS3 dataset, where the number of data channels varies across datasets. In this context, ’filter’ refers to the number of channels in the CNN module, ’kernel’ indicates the size of the convolutional kernel, and ’hidden dim’ represents the size of the hidden layer
ID Layer Type Output Dim Filters Kernel Hidden Dim Activation
1 Input 3000 $\times$ 25 - - - -
2 ST Module 600 $\times$ 125 - - - -
3 Temporal Encoder 300 $\times$ 125 16 31 - ReLU
4 Graph Encoder 300 $\times$ 112 - - 512 ELU
5 Temporal Encoder 150 $\times$ 112 32 17 - ReLU
6 Graph Encoder 150 $\times$ 89 - - 256 ELU
7 Temporal Encoder 75 $\times$ 89 64 9 - ReLU
8 Graph Encoder 75 $\times$ 78 - - 128 ELU
9 Temporal Encoder 15 $\times$ 78 128 5 - ReLU
10 Graph Encoder 15 $\times$ 46 - - 64 ELU
12 Bottom Block 15 $\times$ 46 128 5 64 ReLU/ELU
13 Graph Decoder 15 $\times$ 78 - - 64 ELU
14 Temporal Decoder 75 $\times$ 78 128 5 - ReLU
13 Graph Decoder 75 $\times$ 89 - - 128 ELU
14 Temporal Decoder 150 $\times$ 89 64 9 - ReLU
13 Graph Decoder 150 $\times$ 112 - - 256 ELU
14 Temporal Decoder 300 $\times$ 112 32 17 - ReLU
13 Graph Decoder 300 $\times$ 125 - - 512 ELU
14 Temporal Decoder 600 $\times$ 125 16 31 - ReLU
16 Reshape 3000 $\times$ 25 - - - -
17 Point-wise Classifier 3000 $\times$ 25 5 1 - ReLU
18 Segment-wise Classifier 1 - - - ReLU
Table S. 5: The hyperparameters used for all datasets.
  Parameter Value
Optimizer Adam
Learning rate 0.001
Network dropout rate 0.5
Classifier dropout rate 0.4
weightDecay 0.005
lrFactor 0.5
lrPatience 5
Input dim 3000
Depth 4
Activations {ReLU, ELU}
Conv. filters {16, 32, 64, 128}
GraphConv. hidden layer {1024 512 256 128}
Conv. kernel size {31, 17, 9, 5}
Max-pool kernel size {2, 2, 2, 5}
GraphPool downsample rate {0.9, 0.8, 0.7, 0.6}
Temporal mask coefficient $\delta_{t}$ 0.8
Position mask coefficient $\delta_{p}$ 0.9
Padding same
Batch normalization True
Batch size 64
Training epochs 80
Table S. 6: ST-USleepNet accuracy for each class across all datasets.
Dataset Wake N1 N2 N3 REM
Acc(%) Acc(%) Acc(%) Acc(%) Acc(%)
ISRUC-S1 88.3 56.1 86.1 84.0 84.8
ISRUC-S3 91.3 64.6 87.5 92.0 88.0
MASS-SS3 91.0 58.9 93.2 88.3 92.6
Figure S. 2: Confusion matrix of each dataset.
Figure S. 3: Results of the ablation study conducted across all datasets.
Figure S. 4: We explore the impact of kernel size and hidden dimension on the model’s performance. ST-uSleepNet implements a multi-scale strategy in the temporal prominence network using different convolutional kernels and in the spatial prominence network using different hidden layer dimensions. By comparing various combinations of convolutional kernels and hidden layer dimensions (increasing then decreasing, as in our model’s strategy; consistently using small kernels and hidden layers, referred to as ’small’; consistently using large kernels and hidden layers, referred to as ’large’; decreasing then increasing, referred to as ’reverse’), we found that using decreasing convolutional kernels and hidden layer dimensions in the encoder, and increasing them in the decoder, achieves the most effective multi-scale modeling.
Figure S. 5: We investigate the impact of the number of layers in the U-shaped network on the model’s performance and determined that a four-layer network is the optimal structure.
Figure S. 6: The overview of ST-USleepNet architecture.